OpenStudy (anonymous):

Please help will fan & medal! Which of the following relations is a function? { (1, 2), (1, 3), (1, 4), (1, 5) } { (1, 3), (2, 3), (4, 3), (9, 3) } { (5, 4), (-6, 5), (4, 5), (4, 0) } { (6, -1), (1, 4), (2, 3), (6, 1) }

OpenStudy (anonymous):

B function = 2(x) y-4

OpenStudy (anonymous):

You can only have one input value (x) for every output value (y). So imagine you have a machine. If you put in only one x-value, it gives you only one y-value. Now lets try using this machine with a relation I came up with: {(1,2)(2,3)(2,4)} So if you put in a 1 you get out a 2 If you put in a 2 you get out a 3 If you put in a 2 you get out a 4 Now, this doesn't make sense, because the machine can only give you one input value for every output value. Therefore, if there are repeating x-values, it is not a function, because you will have to graph the same x-value to more than one y-value. And just to make this less confusing in case it is: x/x-value = input/input value y/y-value = output/output value Hope this helps and I'm here if you have questions! :)

OpenStudy (anonymous):

Thank you so much! :) @bishop.courtney2014 & @HorseGirl80

OpenStudy (anonymous):

You're welcome! And thanks for the medal! :D

OpenStudy (anonymous):

I plugged in the wrong number possibly too i think. I thing=k your right @HorseGirl80

OpenStudy (anonymous):

Thanks! This is the first question I've ever answered, so that boosts my confidence! lol ;D